A child is sitting in a chair connected to a rope that passes over a friction less pulley .the child pulls on the loose end of the rope with a force of 250 N .the child's weight is 320 N and the chair weights 160 N .the child is accelerating .find the force that the seat of the chair exerts on the child?


1
Expert's answer
2020-02-28T10:21:04-0500

Weight of the child is 320/10= 32 kg

Force exerted by the child on the chair will be equal to weight of the child - the force applied by the child in pulling the rope

= 320-250 = 70 N